OpenStudy (akohl):

@Abhisar A car of mass 1390kg accelerates on a flat highway from 12 m/s to 17.0 m/s. How much work does the car's engine do on the car?

OpenStudy (abhisar):

\(\huge \sf Work~done = Change~in~Kinetic~energy\)

OpenStudy (akohl):

Your response came out weird. It can't be read

OpenStudy (abhisar):

Work done=Change in Kinetic energy

OpenStudy (akohl):

Right work is force times displacement . So force would be 1390xa and there's no displacements

OpenStudy (akohl):

If I do kinetic energy which velocity do I use

OpenStudy (abhisar):

but you don't know a ! So u can not use W=F\(\times\)Displacement here !

OpenStudy (abhisar):

U have to calculate change in K.E \(\large\sf \frac{1}{2}mv^2-\frac{1}{2}mu^2\)

OpenStudy (akohl):

Whats (\large\sf \frac{1}{2}mv^2-\frac{1}{2}mu^2\) suppose to be

OpenStudy (akohl):

It just shows up as symbols on my computer

OpenStudy (abhisar):

O_o

OpenStudy (abhisar):

http://prntscr.com/44lhyt

OpenStudy (akohl):

I know kinetic is .5*139*v^2. What velocity do I use

OpenStudy (abhisar):

U saw the screenshot ? u r not getting u'll have to use both velocities. You have to calculate change in kinetic energy i.e final kinetic energy -initial kinetic energy.

OpenStudy (akohl):

Oh okay use the equation twice with each and subtract. The answer is the difference thanks

OpenStudy (abhisar):

or u can simple plug in the values in the following equation 1/2*m*(V^2-U^2)

OpenStudy (akohl):

I don't know u

OpenStudy (akohl):

I got 101 kJ

OpenStudy (abhisar):

u mean u don't understand me..ryt ?

OpenStudy (akohl):

I think I got it 101kj. Figured out what u was

OpenStudy (abhisar):

That's correct !

OpenStudy (akohl):

Thanks putting up another question

OpenStudy (abhisar):

\(\color{red}{\huge\bigstar}\huge\text{You're Most Welcome! }\color{red}\bigstar\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)